# TSD-SR: One-Step Diffusion with Target Score Distillation for Real-World Image Super-Resolution

### v2

@@ 3.5 Training Objective
We summarize all the losses we used in our framework.
@@
For the reconstruction loss, we use $LPIPS$ loss:
$\mathcal{L}_{Rec}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)=\mathcal{L}_{LPIPS}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right).$ (7)
@@
$\mathcal{L}_{Stu}=\mathcal{L}_{Rec}+\gamma\mathcal{L}_{Reg},$ (8)
where $\gamma$ is a weighting scalar.
@@ 4.1 Experimental Settings
Implementation Details. All models mentioned are initialized from the Teacher Model (SD3 [10] in our work). We only train vae encoder and denoising network in the Student Model, frozen the vae decoder in order to preserve vae’s prior [24]. We utilize the default prompt for the Student Model, and prompts extracted from HQ for the Teacher and LoRA models when training. We use the AdamW optimizer [30] with a learning rate of 5e-6 for the Student Model and 1e-6 for the LoRA Model, and set rank of LoRA to 64 for both two models. The training process took roughly 48h, utilizing 8 NVIDIA V100 GPUs with a batch size of 16.
@@
DRealSR Ours-1s 26.01 0.7174 0.3092 0.2216 130.83 5.5044 66.66 0.5850 0.7345
OSEDiff-1s 23.94 0.6736 0.3172 0.2363 125.93 6.3822 67.52 0.6187 0.7001
@@
RealSR Ours-1s 23.27 0.6874 0.2807 0.2185 113.31 5.0679 71.28 0.6346 0.7235
OSEDiff-1s 23.72 0.6109 0.3058 0.2138 26.34 5.3903 65.27 0.5838 0.6558
@@ 4.2 Comparison with Existing Methods
Quantitative Comparisons. Tab. 1 show the quantitative comparison of our method with single-step diffusion models on three datasets. Ours achieves the best results on all metrics except PSNR and SSIM. SinSR and AddSR, as distilled versions of previous multi-step super-resolution methods, reduce inference steps but experience a corresponding decrease in performance metrics. OSEDiff introduces the VSD loss from 3D generation tasks into Real-ISR without fully addressing the substantial differences between the two domains. Therefore, the no-reference metrics for image quality are not very satisfied. In contrast, our proposed TSD-SR, tailored for Real-ISR, outperform other single-step models in terms of the vast majority of key metrics.
Tab. 2 show the quantitative comparison with multi-step models. We can draw the following conclusions: (1) TSD-SR demonstrates significant advantages over competing methods in terms of LPIPS, FID, and NIQE metrics. Additionally, it achieves performance that surpasses most multi-step models on DISTS, MUSIQ, and CLIPIQA. (2) DiffBIR, SeeSR, PASD, and AddSR exhibit better performance on the MANIQA metric, which may be attributed to the fact that multi-step models have more denoising iterations to produce rich details. (3) ResShift stands out with the highest PSNR and SSIM scores, while StableSR also shows notable performance regarding DISTS and FID metrics. However, both two models underperform in terms of no-reference metrics.
Finally, we explain the lower PSNR and SSIM metrics in our experiments. Several works [54, 60] have found that these reconstruction metrics are not well-suited for evaluating Real-ISR tasks. When a model recovers better details, it leads to lower reconstruction metrics, indicating a trade-off [3, 67, 32]. Refer to the supplementary material for detailed visual comparisons and analysis.
@@ 4.3 User Study
Datasets Method NIQE $\downarrow$ MUSIQ $\uparrow$ MANIQA $\uparrow$ CLIPIQA $\uparrow$
DRealSR w/o TSM 5.6268 63.90 0.5749 0.6958
w/o DASM 6.0368 63.56 0.5812 0.7123
Full 5.5044 66.66 0.5850 0.7345
RealSR w/o TSM 5.2728 68.56 0.6338 0.6987
w/o DASM 5.5638 69.05 0.6273 0.7031
Full 5.0679 71.28 0.6346 0.7235
4.4 Ablation Study
To validate the effectiveness of TSM loss and DASM, we conducted ablation studies by removing them separately in our experiments. We selected no-reference metrics for comparison, as these are critical for assessing image quality. The results are presented in Tab. 4. We draw the following conclusions: (1) The absence of TSM loss and DASM affects the no-reference metrics NIQE, MUSIQ, MANIQA, and CLIPIQA. (2) The lack of TSM leads to a significant decrease in MUSIQ and CLIPIQA metrics, possibly due to the unreliable directions in VSD causing unrealistic generations. (3) The absence of DASM results in a decline in NIQE, MUSIQ and CLIPIQA, possibly due to the suboptimal optimization of details.
@@ Appendix A Comparison with GAN-based Methods
We compared our method with GAN-based approaches in Tab. 5. While the GAN methods show advantages in full-reference metrics such as PSNR and SSIM, our model outperforms GAN-based methods on all no-reference metrics. Some research has found the limitations of full-reference methods in the field of image super-resolution [60, 54]. Full-reference metrics often assign low scores to visually appealing results because these results produce more details, some of which may not align with the ground truth. However, no-reference metrics assess image quality based on the individual image, without the need to forcibly align with the ground truth. Therefore, in more complex and realistic degradation scenarios, no-reference metrics may be more suitable for evaluating the results of image super-resolution. In Appendix C, we discuss the comparison between full-reference metrics and human preferences, and in the Fig. 9, we present a visual comparison with GAN-based methods. From the visualization, it can be observed that our model achieves better results in terms of texture details compared to GAN methods.
@@ Appendix B More Visual Comparisons
In Figs. 10, 11 and 12, we provide more visual caparisons with other diffusion-based methods. Numerous examples demonstrate the robust restoration capabilities of TSD-SR and the high quality of the restored images.
@@ Appendix C Comparisons of Full-reference Metrics and Human Preference
We present additional comparative experiments to demonstrate the limitations of full-reference metrics in Figure 13. It can be observed that GAN-based methods with advantages in PSNR and SSIM often result in blurred details and textures. In contrast, our results offer better visual effects, yet they do not hold an advantage in terms of PSNR and SSIM metrics. This phenomenon has also been discussed in other research works [3, 67, 32], with the majority of studies suggesting that for models with strong generation and restoration capabilities, these full-reference metrics do not adequately reflect image quality. We anticipate the development of better full-reference metrics in the future to assess advanced Real-ISR methods.
@@ Appendix D Theory of Target Score Matching
The effectiveness of the method can be proven by starting from the KL divergence. We can use a Sticking-the-Landing style gradient by thinking of ${\epsilon}_{\psi}(\mathbf{z}_{t};c_{y},t)$ as a control variate for $\hat{\epsilon}$ For detailed proof, refer to Appendix 4 of DreamFusion [36]. It demonstrates that the gradient of this loss yields the same updates as optimizing the training loss $\mathcal{L}_{\mathrm{MSE}}$ Eq. 10, excluding the Diffusion Jacobian term.
@@ Appendix E Algorithm
Algorithm 1 details our TSD-SR training procedure. We use classifier-free guidance (cfg) for the Teacher Model and the LoRA Model. The cfg value is set to 7.5.

### v3

@@ 3.5 Training Objective
We summarize all the losses that we used in our framework.
@@
For the reconstruction loss, we use $LPIPS$ loss and $MSE$ loss in the latent space:
$\begin{split}&\mathcal{L}_{Rec}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)\\ &=\gamma_{1}\mathcal{L}_{LPIPS}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)+\mathcal{L}_{MSE}\left(\mathbf{z}_{t},\mathbf{\hat{z}}_{t}\right).\end{split}$ (7)
@@
$\mathcal{L}_{Stu}=\mathcal{L}_{Rec}+\gamma_{2}\mathcal{L}_{Reg},$ (8)
where $\gamma_{1}$ and $\gamma_{2}$ are weighting factors. We initialize both $\gamma_{1}$ and $\gamma_{2}$ to 1 at the beginning of training. As optimization progresses, we ramp up $\gamma_{1}$ from 1 to 2 while maintaining $\gamma_{2}$ at its initial value.
@@ 4.1 Experimental Settings
Implementation Details. All models mentioned are initialized from the Teacher Model (SD3 [10] in our work). Similar to OSEDiff [53], we only train the vae encoder and denoising network in the Student Model, frozen the vae decoder in order to preserve vae’s prior [24]. We utilize the default prompt for the Student Model, and prompts are extracted from HQ for the Teacher and LoRA models when training. We use AdamW optimizer [30] with a learning rate of 5e-6 for the Student Model and 1e-6 for the LoRA Model, and set the rank of LoRA to 64 for both two models. During the initial training phase, we incorporate MSE loss in the latent space and exclude DASM to enhance model stability and reduce training time. In later stages, we remove this MSE loss as it leads to over-smoothed results, and we employ DASM to improve image restoration. The training process took roughly 96h, utilizing 8 NVIDIA V100 GPUs with a batch size of 16.
@@
DRealSR Ours-1s 27.77 0.7559 0.2967 0.2136 134.98 5.9131 66.62 0.5874 0.7344
OSEDiff-1s 25.15 0.7341 0.2920 0.2128 123.48 5.6471 69.10 0.6326 0.6687
@@
RealSR Ours-1s 24.81 0.7172 0.2743 0.2104 114.45 5.1298 71.19 0.6347 0.7160
OSEDiff-1s 23.72 0.6109 0.2942 0.1975 26.34 4.7089 67.96 0.6131 0.6681
@@ 4.2 Comparison with Existing Methods
Quantitative Comparisons. Tab. 1 show the quantitative comparison of our method with single-step diffusion models on three datasets. Ours achieves the best results on most evaluation metrics. SinSR and AddSR, as distilled versions of previous multi-step super-resolution methods, reduce inference steps but experience a corresponding decrease in performance metrics. OSEDiff introduces the VSD loss from 3D generation tasks into Real-ISR without fully addressing the substantial differences between the two domains. Therefore, the no-reference metrics for image quality are not very satisfied. In contrast, our proposed TSD-SR, tailored for Real-ISR, outperforms other single-step models in terms of the vast majority of key metrics.
Tab. 2 show the quantitative comparison with multi-step models. We can draw the following conclusions: (1) TSD-SR demonstrates significant advantages over competing methods in terms of LPIPS, DISTS, and NIQE metrics. Additionally, it achieves performance that surpasses most multi-step models on FID, MUSIQ, and CLIPIQA. (2) DiffBIR, SeeSR, PASD, and AddSR exhibit better performance on the MANIQA metric, which may be attributed to the fact that multi-step models have more denoising iterations to produce rich details. (3) ResShift stands out with the highest PSNR and SSIM scores, while StableSR also shows notable performance regarding DISTS and FID metrics. However, both two models underperform in terms of no-reference metrics.
Finally, we explain the lower PSNR and SSIM metrics in our experiments. Several works [55, 61] have found that these reconstruction metrics are not well-suited for evaluating Real-ISR tasks. When a model recovers better details, it leads to lower reconstruction metrics, indicating a trade-off [3, 68, 32]. This phenomenon has also been widely discussed in other related research work [3, 68, 32, 61, 66, 45, 55]. LPIPS [66] is proposed to overcome the limitation that PSNR and SSIM fail to align with human judgments in spatial ambiguities situation. Other DMs-based SR researchers [61, 45] argue that DMs introduce superior pre-trained priors, enabling the restoration of information that traditional methods (from scratch) cannot achieve. However, such capability often leads to a decline in pixel-level metrics, as they prioritize distribution modeling and sampling from learned distributions over strict pixel fidelity. We anticipate the development of better full-reference metrics in the future to assess advanced Real-ISR methods. Refer to the supplementary material for detailed visual comparisons.
@@ 4.3 User Study
4.4 Ablation Study
Effectiveness of TSM and DASM. To validate the effectiveness of TSM loss and DASM, we conduct ablation studies by removing them separately in our experiments. We select LPIPS, DISTS, MUSIQ, MANIQA and CLIPIQA for comparison, as these metrics are critical for image quality assessment. Additionally, we select FID as the metric to evaluate distribution similarity. The results are presented in Tab. 4. We draw the following conclusions: (1) The absence of TSM loss and DASM negatively impacts performance across both reference-based (LPIPS, DISTS) and non-reference-based metrics (MUSIQ, MANIQA and CLIPIQA). The FID metric is also adversely affected, suggesting a decline in distribution fidelity. (2) The lack of TSM leads to a significant decrease in LPIPS, DISTS, MUSIQ, and CLIPIQA metrics, possibly due to the unreliable directions in VSD causing unrealistic generations. (3) The absence of DASM results in a decline in FID, MUSIQ and CLIPIQA, possibly due to the suboptimal optimization of details.
@@ 4.4 Ablation Study
Datasets Method LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ MUSIQ $\uparrow$ MANIQA $\uparrow$ CLIPIQA $\uparrow$
DRealSR w/o TSM 0.3176 0.2327 139.67 63.90 0.5749 0.6958
w/o DASM 0.3097 0.2311 143.97 63.56 0.5812 0.7123
Full 0.2967 0.2136 134.98 66.62 0.5874 0.7344
RealSR w/o TSM 0.2934 0.2397 117.15 68.56 0.6338 0.6987
w/o DASM 0.2873 0.2273 123.51 69.05 0.6273 0.7031
Full 0.2743 0.2104 114.45 71.19 0.6347 0.7160
Base model for fairer comparison. To validate the effectiveness of our method across different versions of SD models, we conduct additional experiments, as shown in Tab. 5, including SD2-base and SD2.1-base models. The performance is evaluated on the DRealSR test dataset [52]. It demonstrates superior performance compared to other one-step SR methods, including OSEDiff [53] and AddSR [55]. Specifically, our SD2-base version model outperforms the single-step AddSR across all perceptual reference metrics and no-reference metrics, particularly excelling in NIQE[64], MUSIQ[22], and CLIPIQA[44], significantly surpassing the performance of one step AddSR. Our SD2.1-base model demonstrates comparable performance to OSEDiff and surpasses it across various metrics, with notable improvements in NIQE and CLIPIQA.
Table 5: Same base model for fairer comparison to validate the effectiveness of TSD-SR.
[…]
AddSR SD2-base 0.3196 0.2242 6.9321 60.85 0.6188
Ours SD2-base 0.3040 0.2234 6.2202 65.14 0.6935
OSEDiff SD2.1-base 0.2968 0.2162 6.4471 64.69 0.6958
Ours SD2.1-base 0.2943 0.2115 5.7934 65.41 0.7109
Parameters $N$ and $s$ in DASM. We provide performance comparisons for different combinations of $N$ and $s$ in Tab. 6. The performance is evaluated on the DRealSR test dataset. $N$ is set to 4, and $s$ is set to 50 in our setting (bold in the table). Larger or smaller $N$ will degrade performance, possibly because it is related to regularization strength. We prefer to use a smaller $N$ because DASM is computationally time-consuming. Therefore, after carefully balancing training duration and model performance, we select $N$=4 as the final value. Small $s$ will have similar performance, but the image quality will be compromised when setting large $s$. Our experimental results indicate that selecting $s$ within the range of 25-75 may yield better performance.
Table 6: Ablation studies for hyperparameter $N$ and $s$.
[…]
2 50 0.3104 0.2327 137.64 64.49 0.5717 0.7118
4 50 0.2967 0.2136 134.98 66.62 0.5874 0.7344
8 50 0.3421 0.2633 151.64 65.73 0.5875 0.7227
4 25 0.3063 0.2201 130.74 66.19 0.5828 0.7269
4 100 0.3176 0.2230 135.54 65.23 0.5782 0.7024
@@ 5 Conclusion and Limitation
6 Acknowledgment
This work was supported by National Natural Science Foundation of China (62293554, U2336212), Zhejiang Provincial Natural Science Foundation of China (LD24F020007), National Key R&D Program of China (SQ2023AAA01005), Ningbo Innovation “Yongjiang 2035” Key Research and Development Programme (2024Z292), and Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).
@@ 6 Acknowledgment
In this Supplementary Material, we provide additional details, including the comparison with GAN-based methods in Appendix A, more visual comparisons in Appendix B, comparisons of full-reference metrics and human preference in Appendix C, theory of Target Score Matching in Appendix D and algorithm in Appendix E. We conduct these additional comparisons and analyses to validate the effectiveness of TSD-SR.
@@ Appendix A Comparison with GAN-based Methods
We compared our method with GAN-based approaches in Tab. 7. While the GAN methods show advantages in full-reference metrics such as PSNR and SSIM, our model outperforms GAN-based methods on all no-reference metrics. Some researchers have found the limitations of PSNR and SSIM in the field of image super-resolution [61, 55]. The effectiveness of PSNR and SSIM in assessing image fidelity in complex degradation scenarios remains debatable, as pixel-level misalignment often arises when restoring severely degraded images. However, no-reference metrics assess image quality based on the individual image, without the need to forcibly align with the ground truth. Therefore, in more complex and realistic degradation scenarios, no-reference metrics may be more suitable for evaluating the results of image super-resolution. In Appendix C, we further provide the visual comparison between full-reference metrics and human preferences, and in Fig. 9, we present a visual comparison with GAN-based methods. From the visualization, it can be observed that our model achieves better results in terms of texture details compared to GAN methods.
@@ Appendix B More Visual Comparisons
In Figs. 10, 11 and 12, we provide more visual comparisons with other diffusion-based methods. Numerous examples demonstrate the robust restoration capabilities of TSD-SR and the high quality of restored images.
@@ Appendix C Comparisons of Full-reference Metrics and Human Preference
We present additional comparative experiments in Figure 13 to demonstrate that PSNR and SSIM may have limitations in assessing image fidelity under complex degradation scenarios. It can be observed that GAN-based methods with higher PSNR and SSIM produce over-smooth or broken textures, raising concerns about their realism and fidelity. While our approach trades off PSNR and SSIM for natural detail restoration, it achieves enhanced realism and broader perceptual acceptance (Our additional user study reveals that 90.28% of participants prefer ours instead of high PSNR and SSIM methods.).
@@ Appendix D Theory of Target Score Matching
The effectiveness of the method can be proven by starting from the KL divergence. We can use a Sticking-the-Landing [39] style gradient by thinking of ${\epsilon}_{\psi}(\mathbf{z}_{t};c_{y},t)$ as a control variate for $\hat{\epsilon}$. For detailed proof, refer to Appendix 4 of DreamFusion [36]. It demonstrates that the gradient of this loss yields the same updates as optimizing the training loss $\mathcal{L}_{\mathrm{MSE}}$ Eq. 10, excluding the Diffusion Jacobian term.
@@ Appendix E Algorithm
Algorithm 1 details our TSD-SR training procedure. We use classifier-free guidance (cfg) for the Teacher Model and the LoRA Model. The cfg weight is set to 7.5.